OpenStudy (anonymous):

If a number is decreased by its reciprocal, the result is 11/30. Find the number.

OpenStudy (anonymous):

(x) - (1/x) = 11/30 Find a common denominator, 30x. Once you do, you can ignore what's in the denominator and solve the numerator. 30x^2 - 30 = 11x 30x^2 - 11x - 30 = 0 30x^2 - 36x + 25x - 30 = 0 6x(5x - 6) + 5(5x - 6) = 0 (6x + 5)(5x - 6) = 0 x = 6/5 and -5/6 There are two solutions.

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

You're welcome :)